# Improving Ab-Initio Cryo-EM Reconstruction with Semi-Amortized Pose Inference

### v1

@@ 4.2 Switching from auto-encoding to auto-decoding
Unlike in early stages when high uncertainty encourages pose exploration, in later stages, as higher frequency details of the structure are resolved, the true pose posterior becomes more concentrated about a single mode. At this point, the gap between the predicted and true posteriors is mainly determined by the error in the pose estimate (predicted mean), prioritizing accuracy over exploration. However, pose inference by a feed-forward network, as a parametric function of the input image, may produce estimates with substantial error, rendering amortization as a barrier to further refinement of the 3D structure. In prior work [11, 42, 43, 44], a similar issue called the amortization gap has been discussed which measures the KL-divergence between the true and predicted variational posteriors.
To address this issue, we adopt a semi-amortized inference scheme [11] comprising two stages. First, the encoder predicts a set of pose candidates using a multi-head architecture. In the second stage, rather than amortized inference, pose parameters are directly optimized for each image using stochastic gradient descent. To initialize poses for the $i$-th image, we choose the one with the lowest reconstruction loss from the set of candidates $\{R_{i,1},\dots,R_{i,M}\}$, namely $R^{*}_{i}=R_{i,s}$ such that
$s=\arg\min_{i}\mathcal{L}_{i}\ .$ (7)
Subsequently, the pose and structure are optimized by coordinate descent using the negative log-likelihood (Eq. 3) as the objective function. Please see Supp. A for more details on pose optimization.
@@ 4.3 Explicit representation as volume decoder
Recent works [10, 5] use coordinate networks [39, 40, 38] to implicitly model the Fourier representation of the 3D structure. Instead, we couple the pose estimation module with an explicit parameterization of the structure in the Fourier domain. The explicit representation is less computationally expensive than an MLP to evaluate and update. Also, this choice is motivated by the fact the implicit decoder needs to be queried multiple times for each image with the multi-head encoder.
@@ 5 Experiments
We use both synthetic and real datasets to compare our semi-amortized ab-initio reconstruction method with state-of-the-art methods, cryoAI [10] that uses amortized inference for the entire reconstruction, and cryoSPARC [8] that performs pose search separately for each image.
@@
We generate synthetic datasets by simulating the image formation process formalized in Sec. 3 using atomic models deposited in the Protein Data Bank (PDB). We compute ground-truth density maps of size $128^{3}$ for a heat shock protein (HSP) [45] ($1.5$), pre-catalytic spliceosome [46] ($4.33$), and SARS-CoV-2 spike protein [47] ($2.13$). Then, $N=50,000$ projections of size $L=128$ are generated by randomly rotating and projecting the density map along the canonical z-axis. Finally, a random CTF is applied in the Fourier space and Gaussian noise is added yielding $\text{SNR}=0.1$.
@@
We use the 80S experimental dataset EMPIAR-10028 [48] containing 105,247 images of length $L=360$ with pixel size $1.34$. We downsample the images to $L=128$ ($3.76$) using cryoSPARC software [8], and randomly split the data into two halves and run the reconstruction methods independently on each. In our method, we make the assumption that particles are centered in the image box, while cryoAI is able to estimate the in-plane translation using a separate encoder. To ensure fair comparison, we use translation parameters estimated by cryoSPARC software and preprocess images to shift particles to the center. To be consistent, we use the same data with cryoAI and disable translation estimation.
@@
During auto-encoding, we use encoders with $M=7$ and $M=15$ heads for reconstruction on synthetic and real datasets, respectively, with Adam [49] to optimize encoder and decoder with learning rates $0.0001$ and $0.05$. Once switched to direct optimization (after $7$ epochs for synthetic and $15$ epochs for real data), we reduce the decoder learning rate to $0.02$ and allocate a new optimizer for pose parameters with learning rate $0.05$. To be consistent with cryoAI, we use a batch size of $64$ and train for the same number of epochs ($20$ for synthetic and $30$ for real data). We use the public cryoAI codebase and run cryoSPARC v4.4.0 [8] with default settings. Methods are implemented in Pytorch [50]. Experiment are run on a single NVIDIA A40 GPU.
Figure 2: Qualitative and quantitative comparison of reconstructions obtained by our proposed semi-amortized method with cryoAI [10] and cryoSPARC [8]. (Left) Final 3D reconstructions on three synthetic datasets and one experimental dataset (EMPIAR-10028) are depicted using ChimeraX [51]. (Right) FSC curves are visualized for quantitative comparison. The red dashed lines show the standard threshold levels of $0.5$ and $0.143$ to report the resolution (in Angstrom) for synthetic and real data, respectively. Our method achieves higher resolution on the Spliceosome and HSP datasets, and it is competitive with the state of the art on the Spike and EMPIAR-10028 datasets.
@@ 5.1 Results
We first qualitatively compare final reconstructions of our semi-amortized approach with cryoAI [10] and cryoSPARC [8] on synthetic and real datasets (Fig. 2, left). Both our method and cryoSPARC capture high-frequency details of the 3D structure on all datasets, whereas reconstructions obtained by cryoAI on the HSP and 80S datasets are inferior. In particular, on HSP, cryoAI gets stuck in local minima as it fails to handle high uncertainty in poses caused by symmetries in this structure.
For quantitative comparison, we visualize the Fourier Shell Correlation (FSC) [52] between the reconstruction and the ground truth (Fig. 2, right). FSC is the gold-standard metric measuring the normalised cross-correlation coefficient between two 3D Fourier volumes along shells of increasing radius. Higher FSC implies more accurate reconstruction. On the Spliceosome, HSP and 80S experimental datasets, our reconstruction outperforms cryoAI. Also, our method outperforms cryoSPARC on Spliceosome and HSP while achieving competitive FSC on Spike and 80S. We use the standard $0.5$ and $0.143$ cutoff thresholds to report the resolution for synthetic and real data, respectively. Our method achieves higher or competitive resolution compared to cryoAI and cryoSPARC on all datasets.
@@
Table 1: Pose accuracy quantified as mean/median errors in units of degrees evaluated on synthetic datasets.
@@
Ours 3.27 / 0.97 1.54 / 0.90 0.68 / 0.61
In Fig. 3, we visualize resolution-time plots, showing that our semi-amortized method gets to high-resolution reconstructions significantly faster than cryoAI. In fact, semi-amortization accelerates the improvement in the resolution and with our explicit structure decoder, we circumvent expensive MLP evaluations and train faster without any degradation to the final reconstruction quality as shown in Fig. 2. Finally, we report the mean and median errors in estimated poses on synthetic datasets in Table 1, showing that our method outperforms cryoSPARC and cryoAI. High errors by cryoAI on HSP dataset shows that it gets stuck in local minima and fails to accurately estimate poses.
@@
Figure 4: Quantitative and qualitative comparison of fully- vs. semi-amortized methods in pose optimization for Spliceosome and Spike datasets (two rows). (Right) The mean of geodesic distances between the predicted pose and ground-truth is visualized at different epochs. By switching from amortized inference to direct optimization, semi-amortized method (blue) enjoys accelerated pose convergence compared to fully-amortized one (red). (Left) Through some examples, the progress of pose inference methods are compared qualitatively. Each plot corresponds to the approximate log posterior for an image, marginalized over in-plane rotations, as view-direction distribution depicted over HEALPix [53] uniform grid on a unit sphere $S^{2}$. After Gnomonic projection to 2D space, the view is limited to a neighborhood around the mode of interest. Black dot is the starting point while blue dots and red dots show poses estimated by fully- and semi-amortized methods, respectively.
@@
Next, we analyze the evolution of pose estimates over the optimization landscape depicted as a heat map (Fig. 4, right). Qualitatively, poses obtained by direct optimization (blue dots) show stable convergence to the optimal point (highlighted area) while those inferred by the encoder (red dots) frequently oscillate. Indeed, the encoder in amortized inference is a globally parameterized function which might be too restrictive, yielding sub-optimal pose parameters. Therefore, poses inferred in an amortized fashion might fail to consistently converge to the optimal point. On other hand, direct optimization is intuitively more flexible as it is performed separately and locally for each image, exhibiting more stable convergence. See Supp. C for more examples.
@@ 5.3 Multi-Modal Pose Posterior
Lastly, we examine how well our multi-head encoder performs vs. cryoAI encoder in terms of handling the uncertainty in the pose on the challenging dataset of HSP [45]. To simplify the visualization, we run our method with $M=4$ heads in this experiment. We first inspect the behavior of encoders for two example cryo-EM images (Fig. 5, left, see Supp. D for more examples). In each row, the approximate posterior distribution over view direction is visualized for both cryoAI and multi-head encoders given the input image and reconstruction. Our multi-head encoder returns a set of plausible candidates, while cryoAI obtains two pose estimates by data augmentation. In both examples, the multi-head encoder identifies the correct mode while cryoAI selects the incorrect one. Note that the pose set predicted by the multi-head encoder contains other posterior modes as well. Intuitively, cryoAI encoder with two pose predictions cannot explore the pose space as much as our multi-head encoder. Thus, the pose posterior in cryoAI remains uncertain and multi-modal during reconstruction. This also contributes to slower convergence to higher-resolution reconstructions as depicted in the top left of Fig. 5. Finally, we visualize the percentage of images assigned to each head by our “winner-takes-all” loss. The relatively even spread of assignments to different heads shows that the multi-head encoder does not suffer hypotheses collapse [30], namely all heads actively participate in the pose prediction. In Supp. B, we investigate how each head specializes in pose prediction for non-overlapping regions of $SO(3)$.
@@ 6 Conclusion, Limitations and Future Work
In this paper, we propose a new semi-amortized approach to ab initio cryo-EM reconstruction. We develop a new multi-head encoder that estimates a set of plausible candidate pose to handle the uncertainty. As the uncertainty is reduced, we transition to an auto-decoding stage where poses are iteratively optimized using SGD for each image. Our results show that our multi-head encoder is able to capture multiple modes of the pose distribution, and our flexible direct optimization enables accelerated convergence of poses and reconstructions. Our method outperforms cryoAI on experimental data and achieves competitive results with cryoSPARC.
Limitations. In this work, we assume that the 3D structure is rigid while they are often flexible and deform within the sample. Also, for simplicity, the particles are assumed to be centered in the frame. As a promising direction for future work, our semi-amortized method with multi-head encoder can be extended to include other latent variables of translation and heterogeneity as well. Moreover, developing a well-defined heuristic to decide on relative length of two stages of auto-encoding and auto-decoding is also an interesting direction to explore in the future.
Societal Impact. Structure determination with Cryo-EM for macromolecular is one of the most exciting areas in structural biology with vast social impact. For instance, cryo-EM tools were used in recovering the structure of the SARS-COV2 spike protein, the discovery of its pre-fusion conformation, and the evaluation of potential medical treatments.
@@
Improving Ab-Initio Cryo-EM Reconstruction with Semi-Amortized Pose Inference
Appendix A Pose Optimization with SGD
During the auto-decoding stage, we alternate between five iterations of pose updates and one iteration of volume update. To update poses, we keep the volume fixed and optimize for the negative log-likelihood (Eq. 3) with respect to the pose parameters. We define the new pose estimate based on the current one as follows:
@@ Appendix A Pose Optimization with SGD
where $R_{\delta}$ is an infinitesimal rotation matrix perturbing the current estimate. We select axis-angle representation to parameterize this perturbation matrix $R_{\delta}$. By a single vector $\boldsymbol{\omega}\in\mathbb{R}^{3}$, one can represent both the axis $||\boldsymbol{\omega}||$ and the angle $0<\frac{\boldsymbol{\omega}}{||\boldsymbol{\omega}||}<\pi$ for any given rotation. Using Rodrigues formula, the perturbation matrix $R_{\delta}(\boldsymbol{\omega})$ can be parameterized as a function of $\boldsymbol{\omega}$. To find the optimal $\boldsymbol{\omega}$, one can initialize it with zero vector, and then use automatic differentiation [54] in pytorch to compute the gradient with respect to $\boldsymbol{\omega}$ and make updates using Adam [49]. However, a naive implementation of the function $R_{\delta}(\boldsymbol{\omega})$ would lead to numerically unstable calculations of the partial derivative $\frac{\partial R_{\delta}}{\partial\boldsymbol{\omega}}$. In fact, there is a singularity at the zero vector, and the partial derivative involves terms that are unstable around the origin. Formally, the derivative of $i$-th column of the rotation matrix $R_{\delta}$ with respect to the vector $\boldsymbol{\omega}$ is [55, 56],
@@
Appendix B Specialization of Encoder Heads

### v2

@@ 4.2 Switching from auto-encoding to auto-decoding
As optimization progresses and fine-grained details of the 3D structure are resolved, the posterior pose variance tends to decrease, with the posterior approaching a unimodal distribution. At this point, the gap between the amortized and variational posterior distributions is mainly determined by the error in the pose estimate (predicted mean), prioritizing accuracy over exploration. However, a feed-forward network, as a globally parameterized function of the input image, may be limited in the accuracy of its predictions, and hence the approximations inherent in amortized inference become significant sources of error in the refinement of the 3D structure. In prior work [11, 42, 43, 44], a related issue, called the amortization gap, is characterized by the KL-divergence between the true and predicted (amortized) variational posteriors.
To address this issue, we adopt a semi-amortized inference scheme [11] comprising two stages. First, the encoder predicts a set of pose candidates using a multi-head architecture. In the second stage, rather than amortized inference, pose parameters are directly optimized for each image using stochastic gradient descent. To initialize poses for the $i$-th image, we choose the one with the lowest reconstruction loss from the set of candidates $\{\phi_{i,1},\,\dots\,,\,\phi_{i,M}\}$, namely $\phi^{*}_{i}=\phi_{i,s}$ such that
$s=\arg\min_{j}\mathcal{L}_{i,j}\ .$ (6)
Subsequently, the pose and structure are optimized by coordinate descent using the negative log-likelihood (Eq. 3) as the objective function. See Supp. A for more details on rotation optimization.
@@ 4.3 Explicit representation as volume decoder
Recent works [10, 5] use coordinate networks [39, 40, 38] to implicitly model the Fourier representation of the 3D structure. Instead, we couple the pose estimation module with an explicit parameterization (3D voxel array) of the structure in the Fourier domain. The explicit representation is less computationally expensive than an MLP to evaluate and update. Also, this choice is motivated by the fact the implicit decoder needs to be queried multiple times for each image with the multi-head encoder.
@@ 5 Experiments
In what follows, we consider both synthetic and real datasets to empirically compare our semi-amortized ab-initio reconstruction method with state-of-the-art methods, cryoAI [10] and cryoDRGN [7] which use amortized inference for reconstruction, and cryoSPARC [8] which performs pose search separately for each particle image.
@@
We generate synthetic datasets by simulating the image formation process formalized in Sec. 3 using atomic models deposited in the Protein Data Bank (PDB). We compute ground-truth density maps of size $128^{3}$ for a heat shock protein (HSP) [45] ($1.5$), pre-catalytic spliceosome [46] ($4.33$), and SARS-CoV-2 spike protein [47] ($2.13$). Then, $N=50,000$ projections of size $L=128$ are generated by randomly rotating and projecting the density map along the canonical z-axis. To simplify the analysis and interpretation of results (e.g. Figs. 4, 5), we consider synthetic particles are centered, and therefore omit the estimation of 2D translation and allowing us to visualize the pose posterior more easily. Finally, a random CTF is applied in the Fourier space and Gaussian noise is added yielding $\text{SNR}=0.1$.
@@
As a widely adopted experimental benchmark, we use the 80S ribosome dataset (EMPIAR-10028 [48]), containing 105,247 particle images of linear box size $L=360$ with pixel size $1.34$. Following cryoAI and cryoDRGN, we downsample the images to $L=128$ ($3.76$) using cryoSPARC software [8], and randomly split the data into two halves and run the reconstruction methods independently on each (to enable Fourier Shell Correlation as a quantitative measure of the resolution of the 3D reconstruction). Unlike synthetic data, particles are not well-centered so it is required to estimate translation parameters as well as rotations. This is achieved by augmenting each head to return translation parameters as well.
@@
During auto-encoding, we use encoders with $M=7$ and $M=15$ heads for reconstruction on synthetic and real datasets, respectively, with Adam [49] to optimize encoder and decoder with learning rates $0.0001$ and $0.05$. Once switched to direct optimization (after $7$ epochs for synthetic and $15$ epochs for real data), we reduce the decoder learning rate to $0.02$ and allocate a new optimizer for pose parameters with learning rate $0.05$. To be consistent with cryoAI, we use a batch size of $64$ and train for the same number of epochs ($20$ for synthetic, $30$ for real data). We use the public cryoAI codebase, run cryoSPARC v4.4.0 [8] with default settings, and run cryoDRGN homogenous ab-initio job. Methods are implemented in Pytorch [50]. Experiments are run on a single NVIDIA A40 GPU.
Figure 2: Qualitative and quantitative comparison of reconstructions obtained by our proposed semi-amortized method, cryoSPIN, with cryoAI [10], cryoDRGN [7] and cryoSPARC [8] We note that cryoAI often becomes stuck in local minima when particles are not centered, so for cryoAI we translate the input images to correct for the spatial offset. (Left) Final 3D reconstructions on three synthetic datasets and one experimental dataset (EMPIAR-10028) are depicted using ChimeraX [51]. (Right) FSC curves are visualized for quantitative comparison. The red dashed lines show the standard threshold levels of $0.5$ and $0.143$ to report the resolution (in Angstrom) for synthetic and real data, respectively. CryoSPIN achieves higher resolution on the Spliceosome and HSP datasets, and it is competitive with the state of the art on the Spike and EMPIAR-10028 datasets.
@@ 5.1 Results
We first qualitatively compare final reconstructions of cryoSPIN with cryoAI [10], cryoDRGN [7], and cryoSPARC [8] on synthetic and real datasets (Fig. 2, left). We observe that cryoAI gets stuck in local minima when evaluated on experimental data (eg, 80S ribosome). In particular, we found that in the presence of unknown planar shift, cryoAI fails to estimate effective translation parameters, but after centering all particles via pre-processing, we could reproduce the cryoAI results. Hence, the provided results for cryoAI in Fig. 2 are obtained after this pre-processing step, whereas our method and others are given off-centered experimental images. Both cryoSPIN and cryoSPARC capture high-frequency details of the 3D structure on all datasets, whereas reconstructions by amortized methods, cryoAI and cryoDRGN, are inferior on the HSP and 80S datasets. In particular, on HSP, cryoAI gets stuck in local minima as it fails to handle high uncertainty in poses caused by symmetries in this structure.
For quantitative comparison, we visualize the Fourier Shell Correlation (FSC) [52] between the reconstruction and the ground truth (Fig. 2, right). FSC is the gold-standard metric measuring the normalised cross-correlation coefficient between two 3D Fourier volumes along shells of increasing radius. Higher FSC implies more accurate reconstruction. On the Spliceosome, HSP and 80S experimental datasets, our reconstruction outperforms amortized methods of cryoAI and cryoDRGN. Also, our method outperforms cryoSPARC on Spliceosome, HSP and 80S, while achieving competitive FSC on Spike. We use the standard $0.5$ and $0.143$ cutoff thresholds to report the resolution for synthetic and real data, respectively. CryoSPIN achieves higher or competitive resolution compared to others on all datasets.
@@
In Fig. 3, we visualize resolution-time plots, showing that cryoSPIN gets to high-resolution reconstructions significantly faster than cryoAI. In fact, semi-amortization in cryoSPIN accelerates the improvement in the resolution and with our explicit structure decoder, we circumvent expensive MLP evaluations and train faster without any degradation to the final reconstruction quality as shown in Fig. 2. In Supp. B, through a detailed comparison with cryoAI, we show cryoSPIN is $\sim 6$x faster and uses $\sim 5$x less memory compared to cryoAI.
Table 1: Mean/median errors of estimated rotations in units of degrees evaluated on synthetic datasets.
@@
CryoSPIN (Ours) 3.27 / 0.97 1.54 / 0.90 0.68 / 0.61
Finally, we report the mean and median errors in estimated poses on synthetic datasets in Table 1, showing that our method outperforms cryoSPARC and cryoAI. High errors by cryoAI on HSP dataset shows that it gets stuck in local minima and fails to accurately estimate poses.
@@
Figure 4: Quantitative and qualitative comparison of fully- vs. semi-amortized methods in pose optimization for Spike (top) and Spliceosome (bottom) datasets. (Left) Mean geodesic distance between the predicted pose and ground-truth is visualized at different epochs. By switching from amortized inference to direct optimization, semi-amortized method (blue) enjoys accelerated pose convergence compared to amortized inference (red). (Right) To visualize pose inference, images depict the approximate log posterior for three particles (marginalized over in-plane rotations) as view-direction distribution over a HEALPix [53] uniform grid on a unit sphere $S^{2}$. After Gnomonic projection to 2D, we show the neighborhood of the mode of interest. Black dots depict starting points. Blue and red dots are pose estimates from fully- and semi-amortized methods.
@@
Next, we analyze the evolution of pose estimates over the optimization landscape depicted as a heat map (Fig. 4, right). Qualitatively, poses obtained by direct optimization (blue dots) show stable convergence to the optimal point (highlighted area) while those inferred by the encoder (red dots) frequently oscillate. Indeed, the encoder in amortized inference is a globally parameterized function which might be too restrictive, yielding sub-optimal pose parameters. Therefore, poses inferred in an amortized fashion might fail to consistently converge to the optimal point. On other hand, direct optimization is intuitively more flexible as it is performed separately and locally for each image, exhibiting more stable convergence. See Supp. D for more examples.
@@ 5.3 Multi-Modal Pose Posterior
Lastly, we examine how well cryoSPIN multi-head encoder performs vs. cryoAI encoder in terms of handling the uncertainty in the pose on the challenging dataset of HSP [45]. To simplify the visualization, we run our method with $M=4$ heads in this experiment. We first inspect the behavior of encoders for two example cryo-EM images (Fig. 5, left, see Supp. E for more examples). In each row, the approximate posterior distribution over view direction is visualized for both cryoAI and multi-head encoders given the input image and reconstruction. Our multi-head encoder returns a set of plausible candidates, while cryoAI obtains two pose estimates by data augmentation. In both examples, the multi-head encoder identifies the correct mode while cryoAI selects the incorrect one. Note that the pose set predicted by the multi-head encoder contains other posterior modes as well. Intuitively, cryoAI encoder with two pose predictions cannot explore the pose space as much as our multi-head encoder. Thus, the pose posterior in cryoAI remains uncertain and multi-modal during reconstruction. This also contributes to slower convergence to higher-resolution reconstructions as depicted in the top left of Fig. 5. Finally, we visualize the percentage of images assigned to each head by our “winner-takes-all” loss. The relatively even spread of assignments to different heads shows that the multi-head encoder does not suffer hypotheses collapse [30], namely all heads actively participate in the pose prediction. In Supp. C, we investigate how each head specializes in pose prediction for non-overlapping regions of $SO(3)$.
@@ 6 Conclusion, Limitations and Future Work
In this paper, we introduce cryoSPIN, a new semi-amortized approach to ab initio cryo-EM reconstruction. We develop a new multi-head encoder that estimates a set of plausible candidate pose to handle the uncertainty. As the uncertainty is reduced, we transition to an auto-decoding stage where poses are iteratively optimized using SGD for each image. Our results show that our multi-head encoder is able to capture multiple modes of the pose distribution, and our flexible direct optimization enables accelerated convergence of poses and reconstructions. CryoSPIN outperforms cryoAI on experimental data and achieves competitive results with cryoSPARC.
Limitations and future work. In this work, we assume that the 3D structure is rigid while many biomolecules are flexible in practice, exhibiting different conformations. Our results show that, even in this relatively constrained homogeneous case, there is a significant gap between amortized inference and our semi-amortized approach. We expect that the benefits of semi-amortized inference will also apply to such heterogeneous cases. In particular, our proposed multi-head encoder should be applicable to the estimation of the multimodal posterior over heterogeneity latent parameters as well as pose. Yet, there are significant challenges in heterogeneous reconstruction, such as extra uncertainty due to heterogeneity and conflation of latent spaces, making it outside the scope of our analysis on semi-amortized inference.
The development of principled criteria for switching from amortized to auto-decoding is also an interesting direction for future research. Our experiments with a range of values show that after a sufficient number of epochs, the reconstruction error becomes insensitive to the switching time. Our intuition is that when switching too early, the pose posterior may have significant uncertainty, and the pose estimates may be too far from the correct basin of attraction to afford robust convergence. Therefore, it is better to switch late than early. However, a well-defined heuristic for the optimal switch time is lacking.
Societal impact. Structure determination with Cryo-EM for macromolecules is one of the most exciting, fundamental areas in structural biology, and is already having with vast social impact. Cryo-EM was used to determine the structure of the SARS-COV2 spike protein, the discovery of its pre-fusion conformation, and the evaluation of medical counter-measures. It compliments advances in methods like alpha-fold for structure prediction, transforming our understanding of the process of life at the scale of the cell, and the design and development of new therapeutics.
@@
CryoSPIN: Improving Ab-Initio Cryo-EM Reconstruction with Semi-Amortized Pose Inference
Appendix A Pose Optimization
For the auto-encoding stage, we follow cryoAI and design the convolutional backbone to output the six-dimensional representation commonly referred to as $S2S2$. To compute the rotation matrix from this representation, the 6D vector is split into two 3D vectors and normalized, denoted as $v_{1},v_{2}\in\mathbb{R}^{3}$. We then compute the cross product between them, $v_{3}=v_{1}\times v_{2}$, yielding a new unit vector. Selecting $v_{1}$ and $v_{3}$ as the first and third columns of the target rotation matrix, we compute $\tilde{v}_{2}=v_{3}\times v_{1}$, which is another unit vector orthogonal to both $v_{1}$ and $v_{3}$. Then, $R=[v_{1},\tilde{v_{2}},v_{3}]$.
Once switched to direct optimization, we change parameterization to axis-angle representation. During auto-decoding, we alternate between five iterations of pose SGD updates and one iteration of volume update. To update poses, we keep the volume fixed and optimize for the negative log-likelihood (Eq. 3) with respect to the pose parameters. We define the new pose estimate based on the current one as follows:
@@ Appendix A Pose Optimization
where $R_{\delta}$ is an infinitesimal rotation matrix perturbing the current estimate. The perturbation matrix $R_{\delta}$ is parameterized by axis-angle representation. By a single vector $\boldsymbol{\omega}\in\mathbb{R}^{3}$, one can represent both the axis $||\boldsymbol{\omega}||$ and the angle $0<\frac{\boldsymbol{\omega}}{||\boldsymbol{\omega}||}<\pi$ for any given rotation. Using Rodrigues formula, the perturbation matrix $R_{\delta}(\boldsymbol{\omega})$ can be parameterized as a function of $\boldsymbol{\omega}$. To find the optimal $\boldsymbol{\omega}$, one can initialize it with zero vector, and then use automatic differentiation [54] in pytorch to compute the gradient with respect to $\boldsymbol{\omega}$ and make updates using Adam [49]. However, a naive implementation of the function $R_{\delta}(\boldsymbol{\omega})$ would lead to numerically unstable calculations of the partial derivative $\frac{\partial R_{\delta}}{\partial\boldsymbol{\omega}}$. In fact, there is a singularity at the zero vector, and the partial derivative involves terms that are unstable around the origin. Formally, the derivative of $i$-th column of the rotation matrix $R_{\delta}$ with respect to the vector $\boldsymbol{\omega}$ is [55, 56],
@@
Appendix B Ablation Study on Decoder
We perform an ablation study on the decoder, detailed in Table 2, as well as a comparison with baselines in terms of reconstruction time per epoch, GPU memory usage, and number of parameters. As our method consists of two stages, we report numbers for each stage separately. In the first stage, the encoder has H=7 heads, while in the second stage, it is replaced with a pose module with size depending on the number of particles ($N$). To facilitate comparison, we include another baseline, cryoAI-explicit, in which the implicit decoder is replaced by an explicit one as in our method. The explicit and implicit decoders have 4.29 and 0.33 million parameters, respectively. Despite a larger decoder, our method is  6x faster and uses  5x less memory compared to cryoAI. Importantly, swapping the implicit decoder with an explicit one (cryoAI-explicit) significantly drops time and memory, indicating that the implicit decoder is a major computational and memory bottleneck. Yet, cryoAI-explicit uses  2x more GPU memory for encoding and is  1.5x slower than our method as it performs early input augmentation and runs the entire encoder twice per image. Our method, by augmenting the encoder head, saves memory and time during pose encoding. Finally, the amortized baseline, which uses an explicit decoder coupled with a multi-head encoder (H=7), uses more memory in decoding (negligible vs implicit decoder) and runs slower than the direct optimization stage of the semi-amortized method.
Table 2: The reconstruction time per epoch, GPU memory, and number of parameters for different methods averaged across three runs. GPU memory is recorded separately for the encoding and decoding modules. Also, two numbers provided for semi-amortized method corresponds to auto-encoding and auto-decoding stages, respectively. In CryoAI-explicit, the implicit decoder of CryoAI is replaced with an explicit decoder. Since we store rotations in 3D representation during direct optimization, the number of parameters of pose module is $N\times 3$ with $N$ denoting number of images in millions.
Model Time (s) GPU Mem. (GB) # Params (M)
Encoding Decoding
Semi-Amortized 99.76, 81.61 1.17, - 2.22, 0.35 13.01, $4.29+N\times 3$
Fully-Amortized 99.75 1.17 2.22 13.01
CryoAI-explicit 142.40 2.32 0.75 9.05
CryoAI 622.39 2.78 14.05 5.09
Appendix C Specialization of Encoder Heads